7. What is the concentration of Fe2+ ions in a saturated solution of FeS at 25°C? The following information (at 25°C) may be useful: Fe (aq) + 2e-→ Fe(s) Fes(s) + 2e. → Fe(s) + S2-(aq) E =-0.44 V E° =-1.01 V (a) Less than 10-20 M b) About 10-20 M c) About 10-10 M d) About 10-M (e) About 0.01 M
Answer #1

Overall cell reaction is:

Fe2+ (aq.) + S2- (aq.) ---------> FeS (s) , E0cell = - 0.44 - ( - 1.01 ) = 0.57 V

For the reverse reaction,

FeS (s) ----------> Fe2+ (aq.) + S2- (aq.), E0cell = - 0.57 V

deltaG0 = - n F E0cell = - 2 * 96500 * ( - 0.57 ) = 110010 J = 110.0 kJ

And

deltaG0 = - R T lnK

lnK = 110.0 / ( - 0.008314 * 298.15)

K = 5.21 * 10-20

Now,

Ksp of FeS = [Fe2+][S2-]

Ksp = S * S

S = sqrt. (Ksp)

S = sqrt. ( 5.21 * 10-20)

S = Solubility of FeS = COnc. of F2+ = conc. of S2- = 2.28 * 10-10

So, (c)